OpenStudy (anonymous):

(X/100) + (Y/66) = (3X+5Y)/Z. What is Z? a. 8 b. 176 c. 330 d. 3630 e. 7260

OpenStudy (anonymous):

Its not K12 @BlossomCake

OpenStudy (anonymous):

\displaystyle J(x) = Li(x) + \sum_{\rho} Li(x^\rho) - \log 2 + \int_{x}^\infty \frac{dt}{t(t^2 - 1)\log t}